OpenStudy (shubhamsrg):

1000 people are standing in a circle, each has a name , 1,2,.. to 1000. 2 is on the right of 1, 3 on the right of 2 .. and so on, and finally, on right of 1000 is again 1. 1 has a sword. He kills 2 and hands over the sword to 3, 3 kills 4 and hands it to 5, and so on. Then 999 kills 1000 and hands it to 1. 1 then kills 3 and gives it to 5 and this process keeps on continuing. So, which number lasts in the end ?

OpenStudy (anonymous):

this is like the open-close door thingy

OpenStudy (anonymous):

so i would say the ones in the end are the ones with numbers that are perfect squares

OpenStudy (shubhamsrg):

what is open-close door thingy? :|

OpenStudy (anonymous):

i am not sure it is the open closed door problem

OpenStudy (anonymous):

for example, if instead of 1000 there are 10, i think the last person standing is 5 (if i read the question correctly)

OpenStudy (texaschic101):

all the even numbers are knocked out on the first go around

OpenStudy (anonymous):

it does look like it. it looks like 1 is open. 2 is close. 3 is open. and so on

OpenStudy (texaschic101):

then every other odd number starting with 3

OpenStudy (anonymous):

here is one version of the door problem, sometimes instead it is lights turning on and off http://www.techinterview.org/post/526370758/100-doors-in-a-row

OpenStudy (shubhamsrg):

5 would have been right for 10 people.

OpenStudy (anonymous):

i think (although i certainly could be wrong) that the difference here is no one comes back to life

OpenStudy (shubhamsrg):

yes, no one comes back to life .-.

OpenStudy (anonymous):

whereas in the doors they open and shut repeatedly

OpenStudy (anonymous):

well being a bonehead i have tried it with 10, 12 and 14 and got 5, 9 and 13 respectively so no pattern has jumped out at me yet

OpenStudy (anonymous):

and 16 leaves me with 1

ganeshie8 (ganeshie8):

looks 1 wil stay alive as long as 'total group count' is even

OpenStudy (anonymous):

i don't think so

OpenStudy (anonymous):

when i did it with 10, 5 was left

OpenStudy (anonymous):

i know for sure 1 is a survivor though

OpenStudy (texaschic101):

after the first round there is nothing left but odd numbers. On the second round, every other odd person is gone, so would you just keep dividing by 2 till you get to the last number ?

ganeshie8 (ganeshie8):

i mean, when we do with 10 1 wil stay alive first round cuz 10 is even 1 will die second round cuz 10/2 = 5

OpenStudy (anonymous):

ooh

OpenStudy (anonymous):

oh wait. that was the answer...my answer is 1...since no one can kill 1...i think

OpenStudy (shubhamsrg):

1 is not right.

OpenStudy (anonymous):

aw

OpenStudy (anonymous):

yeah one always stays around for the first pass, evens go

ganeshie8 (ganeshie8):

with 1000, he wil stay alive till the group count becomes 125

OpenStudy (anonymous):

moving right along... 10 : 5 12 : 9 14 : 13 18 : 1 20 : 9

OpenStudy (anonymous):

22 : 13

OpenStudy (shubhamsrg):

missed 16 ?

OpenStudy (anonymous):

16 : 1

OpenStudy (texaschic101):

There is probably a formula for this

OpenStudy (anonymous):

i think i made a mistake somewhere

OpenStudy (anonymous):

10: 5 12 ; 9 14 : 13 16 ; 1 18 ; 5 20 ; 9 22 : 13

OpenStudy (anonymous):

24 : 17 hmm maybe it is coming ...

ganeshie8 (ganeshie8):

10 to 30 :-

ganeshie8 (ganeshie8):

no pattern really

OpenStudy (shubhamsrg):

I perhaps see a pattern is it 31 : 31 and 32 : 1 ? I am just predicting on the basis of a pattern. Please confirm ?

OpenStudy (anonymous):

looks pretty patterny is 32 : 1 ?

ganeshie8 (ganeshie8):

Correct !!

OpenStudy (anonymous):

oh, what @shubhamsrg said

OpenStudy (shubhamsrg):

hmm for powers of 2, it comes back to 1 !! and increment of 2 takes place after that!

OpenStudy (shubhamsrg):

so 1 at 512 , and 1024 guess we should take case of 1024 and try to revert back ?

OpenStudy (shubhamsrg):

nah, with 512 is easier

ganeshie8 (ganeshie8):

wow !

OpenStudy (anonymous):

start over at powers of 2

OpenStudy (shubhamsrg):

488 terms after 512 so final ans is 488*2 +1 = 977 voila! that is correct! too good.

OpenStudy (anonymous):

man i am late in the game

OpenStudy (anonymous):

of course the real question is not 'what' but "why"

OpenStudy (shubhamsrg):

I had asked the same question 1 year ago. a programming solution was proposed then. thats how i know 977 is the right ans. http://openstudy.com/users/shubhamsrg#/updates/4ff72322e4b01c7be8c9ee6a

OpenStudy (texaschic101):

I agree...its is 977.........check it out http://tierneylab.blogs.nytimes.com/2009/07/09/solutions-and-prizes-for-the-circle-of-fire-problem/

OpenStudy (anonymous):

but why does it start over at powers of 2? knowing that, it is not that hard

OpenStudy (shubhamsrg):

cool !

OpenStudy (shubhamsrg):

maybe because no. people get half after every round. Thus 2 must have some relation to this :|

OpenStudy (loser66):

I construct the series of numbers: 1 round: 1,3,5,7,9,......... with different is 2= 2^1 2nd round: 1, 5,9,13..... with different is 4 =2^2 3rd round : 1,9,17,25,......with different is 8 =2^3 4th round : 1,17,33,49,.........with different is 16 =2^4 so, it may relate to 2 from that??